The denominator of a rational number is greater than its numerator by 7. if 3 is subtracted from the numerator and 2 is added to its denominator,the new number becomes 1/5.Find the number
Answer by lwsshak3(11628)

The denominator of a rational number is greater than its numerator by 7. if 3 is subtracted from the numerator and 2 is added to its denominator,the new number becomes 1/5.Find the number.
**
let n=numerator
n+7=denominator
..
n-3/(n+7)+2=1/5
n-3/n+9=1/5
5n-15=n+9
4n=24
n=6
n+7=13
number: 6/13